How to solve:
3 sin 45 degrees - 4 tan pi/6?
Thanks!

Answer by sudhanshu_kmr(1152)


pi/6 = 180/6 = 30 degree
so, 3sin45 degree - 4 tan 30 degree
= 3 * 1/ root 2 - 4 * 1/root 3